OpenStudy (anonymous):

simplify: cos(2x-y)cosy-sin(2x-y)siny

OpenStudy (anonymous):

cos 2x

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

cos (A+B)= cosAcosB - sinAsinB

OpenStudy (anonymous):

if u plug in A = 2x -y and B=y u get the same form

OpenStudy (anonymous):

woah thanks!

OpenStudy (anonymous):

noprob

OpenStudy (anonymous):

find all solutions in the interval [0,2pie) 3sin(2t-pie)=3

OpenStudy (anonymous):

sin (2t-pi) = 1 2t - pi = pi/2, 5pi/2.... 2t = 3pi/2, 7pi/2 t=3pi/4,7pi/4..and so on..there might be one more solution..check for more

OpenStudy (anonymous):

post new questions in the questions panel from now on